OpenStudy (anonymous):

Ethel has a candy dish on her table. For every 4 gum drops, there are 3 nut clusters. If there are 15 nut clusters in the dish, how many gum drops are in the dish? A. 20 B. 15 C. 28 D. 25

OpenStudy (anonymous):

20

OpenStudy (anonymous):

15/3=x/4 solve for x

OpenStudy (anonymous):

3x=4*15 3x=60 x=60/3 x=20